A bank pays interest continuously at the rate of 6%. How long does it take for a deposit of Q0 to grow in value to 2Q0






1
Expert's answer
2022-02-03T09:33:31-0500

Continuously Compound Interest:


"Q(t)=Q_0e^{rt}"

Then


"e^{rt}=\dfrac{Q}{Q_0}"

"t=\dfrac{\ln(Q/Q_0)}{r}"

Given

"r=0.06, Q=2Q_0"

"t=\dfrac{\ln(2Q_0/Q_0)}{0.06}\approx11.55(years)"